OpenStudy (haleyelizabeth2017):

State the domain and range of the relation y=arccos x. Please no direct answers. Thanks.

OpenStudy (danjs):

first, what is the domain/range for y=cos(x)

OpenStudy (haleyelizabeth2017):

The domain is all real numbers?

OpenStudy (haleyelizabeth2017):

And the range is −1 ≤ y ≤ 1...

OpenStudy (danjs):

yes

OpenStudy (danjs):

the arccos is the inverse function... the range of cos(x) will become the domain of arccos(x)

OpenStudy (haleyelizabeth2017):

Okay

OpenStudy (danjs):

y = cos(x) ------ y = arccos(x) cos(y) = x inverse

OpenStudy (danjs):

domain and range of x = cos(y) domain, x is -1 to 1,

OpenStudy (danjs):

range resulting from that domain cos(y) = -1 at y=pi to cos(y)=1 at y = 0 range 0 to pi

OpenStudy (danjs):

you have to restrict the cosine functions domain , in order to take its inverse cosine is not 1-to-1 function, unless you only take the restricted domain..

OpenStudy (haleyelizabeth2017):

Okay, thanks

OpenStudy (danjs):

inverse is a reflection over line y=x, you can only do that if you take the restricted domain on cos(x) from 0 to pi/2 that is why the inverse cosine is just that one section

OpenStudy (danjs):